OpenStudy (anonymous):

Exponents. What would the number be in exponential notation if the the base is 7 and the standard notation is 5,764,801?

OpenStudy (rulnick):

If you mean scientific notation, it's 5.764801 x 10^6.

sam (.sam.):

7^(x)=5764801 7^(x)=7^(8) x=8

OpenStudy (anonymous):

thanks!!!!!!!!